# Inducing Riesz Bases in $L^{2}$ via Composition Operators

### v2

@@ 1 Introduction
forms a Riesz basis of $L^{2}(\Omega_{2})$ and characterize its dual, see Theorem˜3.1 and Theorem˜3.2, respectively. Given the potential importance of such induced sequences in optimization problems, we restrict our analysis to the case where $h$ is a differentiable mapping. Our results show that all Riesz bases of the form (1.1) are induced precisely by mappings with finite volume distortion; that is, by mappings whose Jacobian determinants are uniformly bounded above and below, see Corollary˜4.1. To illustrate the practical implications of our study, we present a simple numerical example demonstrating how the mapping $h$ can be modelled by an invertible neural network to construct effecitve Riesz bases of the form (1.1).
Our study is motivated by recent computational advances in nuclear-motion theory [18, 24, 26] and condensed matter physics [27, 25], where parameterized and optimized composition operators have been used to construct bases with improved approxinmation properties. While similar problems have been explored in settings such as weighted Hilbert Bergman spaces [12], spaces of holomorphic functions [13], weighted Hardy spaces in the unit disk [6], and Fock spaces [10], a unified perspective for $L^{2}$ spaces remains unexplored. To address this gap, we rely on Singh’s characterization of bounded and invertible composition operators [22], which extends naturally to composition operators between $L^{2}$ spaces on distinct domains.
@@ 3 Inducing Bases via Composition Operators
Since $C_{h}$ is invertible, it follows that $g\mathrel{\mathop{\ordinarycolon}}=C_{h}f$ is a characteristic function if and only if $f$ is a characteristic function, see [21, Corollary 2.2.3]. Denote by $A$ the inverse of $C_{h}$. It follows immediately that $Ag$ is a characteristic function if and only if $f$ is a characteristic function. Since $A$ preserves $L^{2}$ characteristic functions and $\Omega_{1}$ and $\Omega_{2}$ are standard Borel spaces, one can deduce that $A$ is a composition operator, see [21, Theorem 2.1.13]. ∎
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $h$ is injective and there exists $r>0$ such that $g_{h}\geq r$ $\mu$-a.e. on $\Omega_{1}$.
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$ and let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping such that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$. Then $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ is the bi-orthogonal dual of $(C_{h}\gamma_{n})_{n=0}^{\infty}$.
@@
To prove that the sequence $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis it suffices to show that the operator $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection. $C_{h}$ acts boundedly from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Moreover, $(C_{h}\gamma_{n})_{n=0}^{\infty}$ being a Riesz basis implies that $C_{h}$ is a bijection with inverse $C_{h^{-1}}$. This, in turn, implies that $h^{-1}$ is non-singular and $g_{h^{-1}}$ is bounded $\mu$-a.e. on $\Omega_{2}$, see Lemma˜3.1 and Theorem˜2.1. By Theorem˜3.1 it holds that $g_{h^{-1}}$ is bounded away from zero for a.e. $x\in\Omega_{2}$. By combining these results we deduce that $M_{g_{h^{-1}}}$ is a bounded bijection. Therefore, $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection.
To see that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ are bi-orthogonal observe that for any $n,m\in\mathbb{N}$ we have
$\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\gamma_{m}\circ h)\ g_{h^{-1}}\ d\mu$ $=\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\gamma_{m}\circ h)\ dh^{-1}_{\#}\mu$
$=\int_{\Omega_{1}}(\gamma_{n}\circ h\circ h^{-1})\ (\gamma_{m}\circ h\circ h^{-1})\ d\mu$
@@
It remains to show that the pair $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ are dual. In other words, it remains to show that for any $f$ in $L^{2}(\Omega_{2})$ we have
$f=\sum_{n=0}^{\infty}\langle f,M_{g_{h^{-1}}}C_{h}\gamma_{n}\rangle_{\Omega_{2}}\ C_{h}\gamma_{n}.$
Note that $C_{h^{-1}}f\in L^{2}(\Omega_{1})$ for any $f\in L^{2}(\Omega_{2})$. Since $(\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{1})$ it follows that
$C_{h^{-1}}f$ $=\sum_{n=0}^{\infty}\ \langle C_{h^{-1}}f,\gamma_{n}\rangle_{\Omega_{1}}\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{1}}f\circ h^{-1}\ \gamma_{n}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{2}}f\ \gamma_{n}\circ h\ g_{h^{-1}}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \langle f,M_{g_{h^{-1}}}C_{h}\gamma_{n}\rangle_{\Omega_{2}}\ \gamma_{n}.$
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be an orthonormal basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{2})$ if and only if $g_{h}=1$ $\mu$-a.e. on $\Omega_{1}$.
@@
We note that sequences of the form $(W_{h}\gamma_{n})_{n=0}^{\infty}$ appeared in some previous studies [19, 17]. However, these results assumed more regularity on $h$.
@@ 4 Induced Bases via Differentiable Mappings
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be an everywhere differentiable and injective mapping that induces a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Then $(C_{h}\gamma_{n})_{n={0}}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $r\leq\det J_{h}\leq R$ for a.e. $x$ in $\Omega_{2}$ and some $r,R>0$.
@@
$(M_{\det J_{h}}C_{h}\gamma_{n})_{n=0}^{\infty}=(\gamma_{n}\circ h\ \det J_{h})_{n=0}^{\infty}.$
@@ 5 Application: learning a basis with improved approximation properties
Clearly, the functions $f_{1}$ and $f_{2}$ belong to $L^{2}(\mathbb{R})$ and hence they can be approximated in the linear span of Hermite functions $(\gamma_{n})_{n=0}^{N-1}$ for some $N\in\mathbb{N}$. However, Hermite approximations of both functions converge in a suboptimal manner. For $f_{1}$, this is due to the non-smooth behavior of at $x=0$. For $f_{2}$, this is due to its faster decay with respect to the weight $\exp(-x^{2}/2)$ inherent to Hermite functions.
@@
Following our theoretical analysis, $(\gamma_{n}\circ h_{1})_{n=0}^{\infty}$ and $(\gamma_{n}\circ h_{2})_{n=0}^{\infty}$ are bases of $L^{2}(\mathbb{R})$ and the duals are given by $\bigl(\gamma_{n}\circ h_{1}\ (h_{1}^{\prime})\bigr)_{n=0}^{N-1}$, $\bigl(\gamma_{n}\circ h_{2}\ (h_{2}^{\prime})\bigr)_{n=0}^{N-1}$, respectively. Using these results, we computed the expansion coefficients for multiple values of $N$ and compared the errors of the resulting approximations with the errors of approximations obtained by using Hermite functions. Figure˜3 shows the approximation error for different values of $N$. The results show that more accurate approximations can be obtained with our choice of the perturbing functions $h_{1}$ and $h_{2}$. Notably, the approximation of $f_{2}$ in the linear span of $(\gamma_{n}\circ h_{2})_{n=0}^{N-1}$ exhibits a larger convergence rate in comparison to the rate achieved by Hermite approximations. No such effect was observed for approximations of $f_{1}$.
@@
Figure 4: Plotted are the Hermite functions $(\gamma_{n})_{n=0}^{3}$ as solid black lines, along with their perturbations $(\gamma_{n}\circ h_{1})_{n=0}^{3}$ (red lines with triangle markers) and $(\gamma_{n}\circ h_{2})_{n=0}^{3}$ (blue lines with circle markers). The functions corresponding to n=0, 1, 2, 3n=0,\ 1,\ 2,\ 3 are plotted in panels a, b, c, d, respectively.
In summary, we showed in the previous examples that the approximation properties of Hermite functions can be improved by composing them with a proper linear or nonlinear bi-Lipschitz function. This result is, indeed, not restricted to Hermite functions and holds for other bases. Further, we demonstrated that such bi-Lipschitz mappings can be constructed via normalizing flows, i. e., invertible neural networks. While the examples are rather simple, we note that a few recent works have already made important first steps into more complex applications. Specifically, [5] and [18] utilized bi-Lipschitz normalizing flows to optimize non-linear perturbations to orthonormal bases of $L^{2}$ for solving differential equations. [18] reported several orders-of-magnitude increased accuracy and improved convergence rates upon optimizing a perturbation of the underlying basis. Our results serve as a theoretical framework for such applications and open up new avenues for the design and numerical analysis of problem-specific complete sequences.

### v3

# thesisphdthesis

@@ 1 Introduction
forms a Riesz basis of $L^{2}(\Omega_{2})$ and characterize its dual, see Theorem˜3.1 and Theorem˜3.2, respectively. Given the potential importance of such induced sequences in optimization problems, we restrict our analysis to the case where $h$ is a differentiable mapping. Our results show that all Riesz bases of the form (1.1) are induced precisely by mappings with finite volume distortion; that is, by mappings whose Jacobian determinants are bounded away from zero and infinity, see Corollary˜4.1. To illustrate the practical implications of our study, we present a simple numerical example demonstrating how the mapping $h$ can be modelled by an invertible neural network to construct Riesz bases of the form (1.1) with favorable approximation properties.
Our study is motivated by recent computational advances in nuclear-motion theory [18, 24, 26] and condensed matter physics [27, 25], where parameterized and optimized composition operators have been used to construct bases with improved approximation properties. While similar problems have been explored in settings such as weighted Hilbert Bergman spaces [12], spaces of holomorphic functions [13], weighted Hardy spaces in the unit disk [6], and Fock spaces [10], a unified perspective for $L^{2}$ spaces remains unexplored. To address this gap, we rely on Singh’s characterization of bounded and invertible composition operators [22], which extends naturally to composition operators between $L^{2}$ spaces on distinct domains.
@@ 3 Inducing Bases via Composition Operators
Since $C_{h}$ is invertible, it follows that $g\mathrel{\mathop{\ordinarycolon}}=C_{h}f$ is a characteristic function if and only if $f$ is a characteristic function, see [21, Corollary 2.2.3]. Denote by $A$ the inverse of $C_{h}$. It follows immediately that $Ag$ is a characteristic function if and only if $f$ is a characteristic function. Since $A$ preserves $L^{2}$ characteristic functions, and $\Omega_{1}$ and $\Omega_{2}$ are standard Borel spaces, one can deduce that $A$ is a composition operator, see [21, Theorem 2.1.13]. ∎
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a non-singular measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $h$ is injective and there exists $r>0$ such that $g_{h}\geq r$ $\mu$-a.e. on $\Omega_{1}$.
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$ and denote by $(\tilde{\gamma}_{n})_{n=0}^{\infty}$ its dual. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a non-singular measurable mapping such that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$. Then $(M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n})_{n=0}^{\infty}$ is the bi-orthogonal dual of $(C_{h}\gamma_{n})_{n=0}^{\infty}$.
@@
To prove that the sequence $(M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n})_{n=0}^{\infty}$ is a Riesz basis it suffices to show that the operator $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection. $C_{h}$ acts boundedly from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Moreover, $(C_{h}\gamma_{n})_{n=0}^{\infty}$ being a Riesz basis implies that $C_{h}$ is a bijection with inverse $C_{h^{-1}}$. This, in turn, implies that $h^{-1}$ is non-singular and $g_{h^{-1}}$ is bounded $\mu$-a.e. on $\Omega_{2}$, see Lemma˜3.1 and Theorem˜2.1. By Theorem˜3.1 it holds that $g_{h^{-1}}$ is bounded away from zero for a.e. $x\in\Omega_{2}$. By combining these results we deduce that $M_{g_{h^{-1}}}$ is a bounded bijection. Therefore, $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection.
To see that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n})_{n=0}^{\infty}$ are bi-orthogonal observe that for any $n,m\in\mathbb{N}$ we have
$\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\tilde{\gamma}_{m}\circ h)\ g_{h^{-1}}\ d\mu$ $=\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\tilde{\gamma}_{m}\circ h)\ dh^{-1}_{\#}\mu$
$=\int_{\Omega_{1}}(\gamma_{n}\circ h\circ h^{-1})\ (\tilde{\gamma}_{m}\circ h\circ h^{-1})\ d\mu$
@@
It remains to show that the pair $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n})_{n=0}^{\infty}$ are dual. In other words, it remains to show that for any $f$ in $L^{2}(\Omega_{2})$ we have
$f=\sum_{n=0}^{\infty}\langle f,M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n}\rangle_{\Omega_{2}}\ C_{h}\gamma_{n}.$
Note that $C_{h^{-1}}f\in L^{2}(\Omega_{1})$ for any $f\in L^{2}(\Omega_{2})$. Since $(\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{1})$ it follows that
$C_{h^{-1}}f$ $=\sum_{n=0}^{\infty}\ \langle C_{h^{-1}}f,\tilde{\gamma}_{n}\rangle_{\Omega_{1}}\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{1}}f\circ h^{-1}\ \tilde{\gamma}_{n}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{2}}f\ \tilde{\gamma}_{n}\circ h\ g_{h^{-1}}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \langle f,M_{g_{h^{-1}}}C_{h}\tilde{\gamma}_{n}\rangle_{\Omega_{2}}\ \gamma_{n}.$
@@
Let $(\gamma_{n})_{n=0}^{\infty}$ be an orthonormal basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a non-singular measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{2})$ if and only if $g_{h}=1$ $\mu$-a.e. on $\Omega_{1}$.
@@
We note that sequences of the form $(W_{h}\gamma_{n})_{n=0}^{\infty}$ appeared in some previous studies [19]. However, these results assumed more regularity on $h$.
@@ 4 Induced Bases via Differentiable Mappings
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$ and let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be an everywhere differentiable and injective mapping that induces a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Then $(C_{h}\gamma_{n})_{n={0}}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $r\leq\det J_{h}\leq R$ for a.e. $x$ in $\Omega_{2}$ and some $r,R>0$.
@@
$(M_{\det J_{h}}C_{h}\tilde{\gamma}_{n})_{n=0}^{\infty}=(\tilde{\gamma}_{n}\circ h\ \det J_{h})_{n=0}^{\infty},$
where $(\tilde{\gamma}_{n})_{n=0}^{\infty}$ denotes the dual of $(\gamma_{n})_{n=0}^{\infty}$.
@@ 5 Application: learning a basis with improved approximation properties
Clearly, the functions $f_{1}$ and $f_{2}$ belong to $L^{2}(\mathbb{R})$ and hence they can be approximated in the linear span of Hermite functions $(\gamma_{n})_{n=0}^{N-1}$ for some $N\in\mathbb{N}$. However, Hermite approximations of both functions converge in a suboptimal manner. For $f_{1}$, this is due to the non-smooth behavior at $x=0$. For $f_{2}$, this is due to its faster decay with respect to the weight $\exp(-x^{2}/2)$ inherent to Hermite functions.
@@
Following our theoretical analysis, $(\gamma_{n}\circ h_{1})_{n=0}^{\infty}$ and $(\gamma_{n}\circ h_{2})_{n=0}^{\infty}$ are bases of $L^{2}(\mathbb{R})$ and the duals are given by $\bigl(\gamma_{n}\circ h_{1}\ (h_{1}^{\prime})\bigr)_{n=0}^{N-1}$, $\bigl(\gamma_{n}\circ h_{2}\ (h_{2}^{\prime})\bigr)_{n=0}^{N-1}$, respectively. Using these results, we computed the expansion coefficients for multiple values of $N$ and compared the errors of the resulting approximations with the errors of approximations obtained by using Hermite functions. Figure˜3 shows the approximation error for different values of $N$. The results show that more accurate approximations can be obtained with our choice of the perturbing functions $h_{1}$ and $h_{2}$. Notably, the approximation of $f_{2}$ in the linear span of $(\gamma_{n}\circ h_{2})_{n=0}^{N-1}$ exhibits both higher accuracy and a faster convergence rate compared to the standard Hermite approximations. In contrast, for $f_{1}$, the convergence rates of the Hermite and perturbed Hermite approximations are comparable.
@@
Figure 4: Plotted are the Hermite functions $(\gamma_{n})_{n=0}^{3}$ (solid black lines), along with their perturbations $(\gamma_{n}\circ h_{1})_{n=0}^{3}$ (red lines with triangle markers) and $(\gamma_{n}\circ h_{2})_{n=0}^{3}$ (blue lines with circle markers). The functions corresponding to n=0, 1, 2, 3n=0,\ 1,\ 2,\ 3 are plotted in panels a, b, c, d, respectively.
In summary, we showed in the previous examples that the approximation properties of Hermite functions can be improved by composing them with a proper linear or nonlinear function of bounded volume distortion. This result is, indeed, not restricted to Hermite functions and holds for other bases. Further, we demonstrated that the inducing mappings can be constructed via normalizing flows, i. e., invertible neural networks. While the examples are rather simple, we note that a few recent works have already made important first steps into more complex applications. Specifically, [5] and [18] utilized bi-Lipschitz normalizing flows to optimize non-linear perturbations to orthonormal bases of $L^{2}$ for solving differential equations. [18] reported several orders-of-magnitude increased accuracy and improved convergence rates upon optimizing a perturbation of the underlying basis. Our results serve as a theoretical framework for such applications and open up new avenues for the design and numerical analysis of problem-specific complete sequences.